OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify sqrt(147) and provide the simplified form. You can show a square root by writing "sqrt()". For example, can be written as "sqrt(2)".

OpenStudy (anonymous):

since 147 is a product of 49 and 3, sqrt(149) can be written as sqrt(49 * 3). This is same as saying sqrt(147) is a product of sqrt(49) and sqrt(3). Simplifying this, sqrt(147) is a product of 7 and sqrt(3). This is 7 times sqrt(3). \[ \sqrt{147} = \sqrt{49 \times 3} = \sqrt{49} \times \sqrt{3} = 7 \sqrt{3} \]